# Datalog-Expressibility for Monadic and Guarded Second-Order Logic

### v2

@@ 3 Homomorphism-Closed GSO
In this section, we prove that the class of finite models of a GSO sentence is a finite union of CSPs of $\omega$-categorical structures whenever its complement is closed under homomorphisms. In particular, every CSP in GSO (and therefore every CSP in MSO) is the CSP of an $\omega$-categorical structure. CSPs that can be formulated as the CSP of an $\omega$-categorical structure have been characterised [15]; this characterisation will be recalled next.
@@ 3.1 CSPs for Countably Categorical Structures
By the theorem of Ryll-Nardzewski, a countable structure $\mathfrak{B}$ is $\omega$-categorical if and only if for every $n\in{\mathbb{N}}$ there are finitely many orbits of the componentwise action of the automorphism group $\operatorname{Aut}(\mathfrak{B})$ of $\mathfrak{B}$ on $B^{n}$ (see, e.g., [34]). There is a known condition that characterises classes of structures that are CSPs of $\omega$-categorical structures. Let $\tau$ be a relational signature and let ${\mathcal{C}}$ be a class of finite $\tau$-structures. Let $\Lambda_{n}$ be the class of primitive positive $\tau$-formulas with free variables $x_{1},\dots,x_{n}$ whose canonical database is in ${\mathcal{C}}$. We define $\sim^{\mathcal{C}}_{n}$ to be the equivalence relation on $\Lambda_{n}$ such that $\phi_{1}\sim^{\mathcal{C}}_{n}\phi_{2}$ holds if for all primitive positive $\tau$-formulas $\psi(x_{1},\dots,x_{n})$ we have that $\phi_{1}(x_{1},\dots,x_{n})\wedge\psi(x_{1},\dots,x_{n})$ is satisfiable in a structure from ${\mathcal{C}}$ if and only if $\phi_{2}(x_{1},\dots,x_{n})\wedge\psi(x_{1},\dots,x_{n})$ is satisfiable in a structure from ${\mathcal{C}}$ [^3] A closely related concept was considered for graph classes in parametrised complexity [27, Definition 6.7.5]; we thank Antoine Mottet for pointing this out to us. The index of an equivalence relation is the number of its equivalence classes. If $\sim$ is an equivalence relation on a set $A$, and $a\in A$, then $[a]_{\sim}$ denotes the equivalence class of $\sim$; if the reference to $\sim$ is clear, we omit $\sim$ in this notation.
@@
The structure $\mathfrak{B}_{1}\coloneqq({\mathbb{Z}};<)$ is not $\omega$-categorical. However, $\sim_{n}^{\operatorname{CSP}(\mathfrak{B}_{1})}$ has finite index for all $n$, and indeed $\operatorname{CSP}({\mathbb{Z}};<)=\operatorname{CSP}({\mathbb{Q}};<)$ and $({\mathbb{Q}};<)$ is $\omega$-categorical. On the other hand, for $\mathfrak{B}_{2}\coloneqq({\mathbb{Z}};\text{Succ})$ we have that the index $\sim_{2}^{\operatorname{CSP}(\mathfrak{B}_{2})}$ is infinite, and it follows that there is no $\omega$-categorical structure $\mathfrak{B}$ such that $\operatorname{CSP}(\mathfrak{B}_{2})=\operatorname{CSP}(\mathfrak{B})$; see [8].
A rich source of examples of $\omega$-categorical structures are structures with finite relational signature that are homogeneous, i.e., every isomorphism between finite substructures can be extended to an automorphism. There are uncountably many countable homogeneous digraphs with pairwise distinct CSP, and it follows that there are homogeneous digraphs with undecidable CSPs. The age of a structure $\mathfrak{B}$ is the class of all finite structures that embed into $\mathfrak{B}$. A structure $\mathfrak{B}$ is called finitely bounded if there exists a finite set $\mathcal{F}$ of finite structures such that a finite structure $\mathfrak{A}$ belongs to the age of $\mathfrak{B}$ if and only if no structure in ${\mathcal{F}}$ embeds into $\mathfrak{A}$.
It is well-known that if a structure is $\omega$-categorical, then all of its reducts are $\omega$-categorical as well [34]. Moreover, it is easy to see that the CSP of reducts of finitely bounded structures is in NP. It has been conjectured that the CSP of reducts of finitely bounded homogeneous structures is in P or NP-complete [17]. This conjecture generalises the finite-domain complexity dichotomy that was conjectured by Feder and Vardi [29] and proved by Bulatov [19] and by Zhuk [45]. However, the generalisation of this result to finitely bounded homogeneous structures is wide open [40].
@@ 3.2 Quantifier Rank
In order to construct $\omega$-categorial structures for a given CSP in GSO, we need to verify the condition given in Theorem 3.1. In this context, it will be convenient to work with signatures that also contain constant symbols. The quantifier rank of a second-order $\tau$-formula $\Phi$ is the maximal number of nested (first-order or second-order) quantifiers in $\Phi$; for this definition, we view $\Phi$ as a second-order sentence with guarded semantics, just as in [6]. If $\mathfrak{A}$ and $\mathfrak{B}$ are $\tau$-structures and $q\in{\mathbb{N}}$ we write $\mathfrak{A}\equiv^{\operatorname{GSO}}_{q}\mathfrak{B}$ if $\mathfrak{A}$ and $\mathfrak{B}$ satisfy the same GSO $\tau$-sentences of quantifier rank at most $q$.
@@
If $\mathfrak{A}$ is a $\tau$-structure and $\bar{a}$ is a $k$-tuple of elements of $A$, then we write $(\mathfrak{A},\bar{a})$ for a $\tau\cup\{c_{1},\dots,c_{k}\}$-structure expanding $\mathfrak{A}$ where $c_{1},\dots,c_{k}$ denote fresh constant symbols being mapped to the corresponding entries of $\bar{a}$. If $\mathfrak{A}$ and $\mathfrak{B}$ are $\tau$-structures and $\bar{a}\in A^{k}$, $\bar{b}\in B^{k}$, and when writing $(\mathfrak{A},\bar{a})\equiv^{\operatorname{GSO}}_{q}(\mathfrak{B},\bar{b})$ we implicitly assume that we have chosen the same constant symbols for $\bar{a}$ and for $\bar{b}$.
@@
In the following, $\tau$ denotes a finite relational signature.
@@
We write $\mathfrak{A}\uplus\mathfrak{B}$ for the structure with domain $A\cup B$ such that $R^{\mathfrak{A}\uplus\mathfrak{B}}\coloneqq R^{\mathfrak{A}}\cup R^{\mathfrak{B}}$ for each relation symbol $R\in\tau$ and $c^{\mathfrak{A}\uplus\mathfrak{B}}=c^{\mathfrak{A}}=c^{\mathfrak{B}}$ for each constant symbol $c\in\rho$.
The following theorem in the special case of $n=0$ is Proposition 4.1 in [6].
@@
Proof 3.9.
Let $\tau$ be the signature of ${\mathcal{C}}$, and let $\Phi$ be a GSO $\tau$-formula with quantifierrank $q$ such that ${\mathcal{C}}=\llbracket\Phi\rrbracket$. By Theorem 3.1 it suffices to show that the equivalence relation $\sim^{\mathcal{C}}_{n}$ has finite index for every $n\in{\mathbb{N}}$. Let $\rho\coloneqq\{c_{1},\dots,c_{n}\}$ be a set of new constant symbols. By Lemma 3.3, there exists an $m\in{\mathbb{N}}$ such that $\equiv_{q}^{\operatorname{GSO}}$ has $m$ equivalence classes on $(\tau\cup\rho)$-structures. If $\phi(x_{1},\dots,x_{n})$ is a primitive positive $\tau$-formula, then define ${\mathfrak{S}}_{\phi}$ to be the $(\tau\cup\rho)$-structure whose elements are the equivalence classes of the smallest equivalence relation on the variables of $\phi$ that contains all pairs $x,y$ such that $\phi$ contains the conjunct $x=y$, and such that $(C_{1},\dots,C_{n})\in R^{\mathfrak{S}}$ for $R\in\tau$ if and only if there are $y_{1}\in C_{1},\dots,y_{n}\in C_{2}$ such that $R(y_{1},\dots,y_{n})$ is a conjunct of $\phi$; finally, we set $c_{i}^{\mathfrak{S}_{\phi}}\coloneqq[x_{i}]$ for all $i\in\{1,\dots,n\}$.
@@
It is easy to see that $\llbracket\Phi\rrbracket$ is closed under disjoint unions and that its complement is closed under homomorphisms. Corollary 3.8 implies that there exists a countable $\omega$-categorical structure with $\operatorname{CSP}(\mathfrak{B})=\llbracket\Phi\rrbracket$. We claim that this statement cannot be deduced from the theorem of Cherlin, Shelah, and Shi [23]. To see this, first note that the following graphs $W_{i}$, for $i\geq 2$, satisfy the mentioned extension property: the vertex set of $W_{i}$ is $\{0,1,\dots,i\}$, there is an edge from $p$ to $q$, for $p,q\in\{0,\dots,i-1\}$, if $p-q=1\mod i$, and additionally we have the edges $(p,i)$ and $(i,p)$ for every $p\in\{0,\dots,i-1\}$. See Figure 2 for an illustration of $W_{6}$. Note that for every $i\geq 2$ the graph $W_{i}$ is a core, and that they are pairwise homomorphically incomparable. Moreover, every subgraph of $W_{i}$ does not satisfy the extension property.
@@ 3.3 Finite Unions of CSPs
In this section we prove that every class in GSO whose complement is closed under homomorphisms is a finite union of CSPs (Lemma 3.17); the statement announced at the beginning of Section 3 then follows (Corollary 3.19). Throughout this section, let $\tau$ be a relational signature and let ${\mathcal{C}}$ be a non-empty class of finite $\tau$-structures whose complement is closed under homomorphisms. In particular, ${\mathcal{C}}$ contains the structure $\mathfrak{I}$ with only one element where all relations are empty.
@@
If $\mathfrak{A}\in{\mathcal{C}}$, then we write $[\mathfrak{A}]$ for the equivalence class of $\mathfrak{A}$ with respect to $\sim$. The following observations are immediate consequences from the definitions:
@@
    each $\sim$-equivalence class is closed under homomorphic equivalence.
@@
Let $\mathfrak{A}\in{\mathcal{C}}$ and let $\mathcal{D}$ be the smallest subclass of $\mathcal{C}$ that contains $[\mathfrak{A}]$ and whose complement is closed under homomorphisms. Then
@@
[…]
@@
To prove the second statement, first observe that the statement is clear if $\mathfrak{A}\in[\mathfrak{I}]$, since the complement of $[\mathfrak{I}]$ is closed under homomorphisms. The statement therefore follows from the assumption that $\sim$ has more than one equivalence class. Otherwise, if $\mathfrak{A}\notin[\mathfrak{I}]$, then there exists a structure $\mathfrak{B}\in{\mathcal{C}}$ such that $\mathfrak{A}\uplus\mathfrak{B}\notin{\mathcal{C}}$. Then $\mathfrak{B}\in\mathcal{C}\setminus{\mathcal{D}}$ can be shown indirectly as follows: otherwise $\mathfrak{B}$ would have a homomorphism to a structure $\mathfrak{A}^{\prime}\in[\mathfrak{A}]$. Since $\mathfrak{B}\uplus\mathfrak{A}^{\prime}$ is homomorphically equivalent to $\mathfrak{A}^{\prime}$, we have $\mathfrak{B}\uplus\mathfrak{A}^{\prime}\sim\mathfrak{A}^{\prime}\sim\mathfrak{A}$ and in particular $\mathfrak{B}\uplus\mathfrak{A}^{\prime}\in{\mathcal{C}}$. But $\mathfrak{B}\uplus\mathfrak{A}^{\prime}\in{\mathcal{C}}$ if and only if $\mathfrak{B}\uplus\mathfrak{A}\in{\mathcal{C}}$ since $\mathfrak{A}\sim\mathfrak{A}^{\prime}$. This is in contradiction to our assumption on $\mathfrak{B}$.
@@
For the remainder of the section we fix a GSO $\tau$-sentence $\Phi$ of quantifier rank $q$. Recall that Lemma 3.3 asserts that the equivalence relation $\equiv^{\operatorname{GSO}}_{q}$ on the class of finite $\tau$-structures has finitely many equivalence classes $\mathcal{C}_{1},\dots,\mathcal{C}_{m}$, and that each of the equivalence classes $\mathcal{C}_{i}$ can be defined by a single GSO $\tau$-sentence $\Psi_{i}$ with quantifier rank $q$; we write $T^{\tau}_{q}\coloneqq\{\Psi_{1},\dots,\Psi_{m}\}$ for this set of GSO sentences. Let $J\subseteq\{1,\ldots,m\}$ be such that $\{\Psi_{j}\in T^{\tau}_{q}\mid j\in J\}$ is exactly the set of all sentences in $T^{\tau}_{q}$ that imply $\Phi$. Then $|J|$ is called the degree of $\Phi$. It is easy to see that the degree of $\Phi$ is exactly the index of $\equiv^{\operatorname{GSO}}_{q}$ restricted to $\llbracket\Phi\rrbracket$. Let $\sim$ be the equivalence relation defined in the beginning of this section for the class ${\mathcal{C}}\coloneqq\llbracket\Phi\rrbracket$.
@@
For every $\sim$-class $\mathcal{D}$ there exists $I\subseteq\{1,\ldots,m\}$ such that ${\mathcal{D}}=\bigcup_{i\in I}\llbracket\Psi_{i}\rrbracket$.
Proof 3.15.

### v3

@@ 3 Homomorphism-Closed GSO
In this section, we prove that the class of finite models of a GSO sentence is a finite union of CSPs of $\omega$-categorical structures whenever its complement is closed under homomorphisms. In particular, every CSP in GSO (and therefore every CSP in MSO) is the CSP of an $\omega$-categorical structure $\mathfrak{B}$. This result will be essential for our proof of the existence of canonical Datalog programs for GSO in Section 4. In our proof, we use an equivalent characterisation of the CSPs that can be formulated as the CSP of an $\omega$-categorical structure from [17]; this characterisation will be recalled in Section 3.1.
Remark 3.1.
This remark is addressed to readers familiar with the universal-algebraic approach to CSPs (see [10] for an introduction), and can be skipped. Recall that whether the CSP of an $\omega$-categorical structure $\mathfrak{B}$ is in Datalog only depends on the polymorphism clone of $\mathfrak{B}$ (see [10, Corollary 8.3.4] for an even stronger statement). Therefore, our result opens up the repertoire of methods from universal algebra for further studies of the intersection of GSO and Datalog.
@@ 3.1 CSPs for Countably Categorical Structures
Given permutation group $G$ on a set $B$, an orbit of $G$ is a set of the form $\{g(b)\mid g\in G\}$ for some $b\in B$. By the theorem of Ryll-Nardzewski, a countable structure $\mathfrak{B}$ is $\omega$-categorical if and only if for every $n\in{\mathbb{N}}$ there are finitely many orbits of the componentwise action of the automorphism group $\operatorname{Aut}(\mathfrak{B})$ of $\mathfrak{B}$ on $B^{n}$ (see, e.g., [37]). It follows that if a structure is $\omega$-categorical, then all of its reducts are $\omega$-categorical as well [37].
There is a known condition that characterises classes of structures that are CSPs of $\omega$-categorical structures. Let $\tau$ be a relational signature and let $\mathcal{C}$} be a class of finite $\tau$-structures. Let $\Lambda_{n}$ be the class of primitive positive $\tau$-formulas with free variables $x_{1},\dots,x_{n}$ whose canonical database is in $\mathcal{C}$}. We define $\sim^{\mathcal{C}}_{n}$ to be the equivalence relation on $\Lambda_{n}$ such that $\phi_{1}\sim^{\mathcal{C}}_{n}\phi_{2}$ holds if for all primitive positive $\tau$-formulas $\psi(x_{1},\dots,x_{n})$ we have that $\phi_{1}(x_{1},\dots,x_{n})\wedge\psi(x_{1},\dots,x_{n})$ is satisfiable in a structure from $\mathcal{C}$} if and only if $\phi_{2}(x_{1},\dots,x_{n})\wedge\psi(x_{1},\dots,x_{n})$ is satisfiable in a structure from $\mathcal{C}$}. [^4] A closely related concept was considered for graph classes in parametrised complexity [30, Definition 6.7.5]; we thank Antoine Mottet for pointing this out to us. The index of an equivalence relation is the number of its equivalence classes. If $\sim$ is an equivalence relation on a set $A$, and $a\in A$, then $a]_{\sim}$ denotes the equivalence class of $a$; if the reference to $\sim$ is clear, we omit $\sim$ in this notation.
@@
The structure $\mathfrak{B}_{1}\coloneqq({\mathbb{Z}};<)$ is not $\omega$-categorical. However, $\sim_{n}^{\operatorname{CSP}(\mathfrak{B}_{1})}$ has finite index for all $n$, and indeed $\operatorname{CSP}({\mathbb{Z}};<)=\operatorname{CSP}({\mathbb{Q}};<)$ and $({\mathbb{Q}};<)$ is $\omega$-categorical. On the other hand, for $\mathfrak{B}_{2}\coloneqq({\mathbb{Z}};\text{Succ})$ we have that the index $\sim_{2}^{\operatorname{CSP}(\mathfrak{B}_{2})}$ is infinite, and it follows that there is no $\omega$-categorical structure $\mathfrak{B}$ such that $\operatorname{CSP}(\mathfrak{B}_{2})=\operatorname{CSP}(\mathfrak{B})$; see [10]. $\triangle$
Remark 3.4.
Since we make essential use of Theorem 3.2, it might be helpful for the reader to have the following idea about its proof. We use the equivalence relation $\sim_{n}^{\mathcal{C}}$ to expand the structures in $\mathcal{C}$} by new relations; for the resulting class we can use Fraïssé’s construction of homogeneous structures (Definition 6.1). The assumption that $\sim_{n}^{\mathcal{C}}$ has finite index is the reason that the resulting structure is also $\omega$-categorical. By taking a reduct to the original signature, we might lose homogeneity, but keep $\omega$-categoricity, and obtain the desired structure $\mathfrak{B}$. $\triangle$
@@ 3.2 Quantifier Rank
In order to construct $\omega$-categorial structures for a given CSP in GSO, we need to verify the condition given in Theorem 3.2. In this context, it will be convenient to work with signatures that also contain constant symbols. The quantifier rank of a second-order $\tau$-formula $\Phi$ is the maximal number of nested (first-order or second-order) quantifiers in $\Phi$; for this definition, we view $\Phi$ as a second-order sentence with guarded semantics, just as in [7]. If $\mathfrak{A}$ and $\mathfrak{B}$ are $\tau$-structures and $q\in{\mathbb{N}}$ we write $\mathfrak{A}\equiv^{\operatorname{GSO}}_{q}\mathfrak{B}$ if $\mathfrak{A}$ and $\mathfrak{B}$ satisfy the same GSO $\tau$-sentences of quantifier rank at most $q$. The following lemma is a natural extension of a classical result of Fräissé and Hintikka on first-order logic (see, e.g., [37]).
@@
If $\mathfrak{A}$ is a $\tau$-structure and $\bar{a}$ is a $k$-tuple of elements of $A$, then we write $(\mathfrak{A},\bar{a})$ for a $\tau\cup\{c_{1},\dots,c_{k}\}$-structure expanding $\mathfrak{A}$ where $c_{1},\dots,c_{k}$ denote fresh constant symbols being mapped to the corresponding entries of $\bar{a}$. If $\mathfrak{A}$ and $\mathfrak{B}$ are $\tau$-structures and $\bar{a}\in A^{k}$, $\bar{b}\in B^{k}$, and when writing $(\mathfrak{A},\bar{a})\equiv^{\operatorname{GSO}}_{q}(\mathfrak{B},\bar{b})$, we implicitly assume that we have chosen the same constant symbols for $\bar{a}$ and for $\bar{b}$. The following lemma is a natural extension of classical results of Ehrenfeucht and Fräissé on first-order logic (see, e.g., [37]).
@@
In the following, $\tau$ denotes a finite relational signature. Theorem 3.8 below states that the $\equiv_{q}^{\text{GSO}}$-equivalence class of the union of two structures which might overlap on some elements that are named by constants can be recovered from the equivalence classes of its components.
@@
Given a $(\mathfrak{A},\mathfrak{B})\in\mathcal{D}_{n}$, we write $\mathfrak{A}\uplus\mathfrak{B}$ for the structure with domain $A\cup B$ such that $R^{\mathfrak{A}\uplus\mathfrak{B}}\coloneqq R^{\mathfrak{A}}\cup R^{\mathfrak{B}}$ for each relation symbol $R\in\tau$ and $c^{\mathfrak{A}\uplus\mathfrak{B}}=c^{\mathfrak{A}}=c^{\mathfrak{B}}$ for each constant symbol $c\in\rho$.
The special case of the following theorem for $n=0$ (i.e., for disjoint unions with no overlap) is Proposition 4.2 in [8].
@@
Proof 3.11.
Let $\tau$ be the signature of $\mathcal{C}$}, and let $\Phi$ be a GSO $\tau$-formula with quantifier rank $q$ such that ${\mathcal{C}}=\llbracket\Phi\rrbracket$. By Theorem 3.2 it suffices to show that the equivalence relation $\sim^{\mathcal{C}}_{n}$ has finite index for every $n\in{\mathbb{N}}$. Let $\rho\coloneqq\{c_{1},\dots,c_{n}\}$ be a set of new constant symbols. By Lemma 3.5, there exists an $m\in{\mathbb{N}}$ such that $\equiv_{q}^{\operatorname{GSO}}$ has $m$ equivalence classes on $(\tau\cup\rho)$-structures. If $\phi(x_{1},\dots,x_{n})$ is a primitive positive $\tau$-formula, then define ${\mathfrak{S}}_{\phi}$ to be the following $(\tau\cup\rho)$-structure:
  • the elements are the equivalence classes of the smallest equivalence relation on the variables of $\phi$ that contains all pairs $x,y$ such that $\phi$ contains the conjunct $x=y$, and
  • $(C_{1},\dots,C_{n})\in R^{\mathfrak{S}}$ for $R\in\tau$ if and only if there are $y_{1}\in C_{1},\dots,y_{n}\in C_{n}$ such that $R(y_{1},\dots,y_{n})$ is a conjunct of $\phi$;
  • finally, we set $c_{i}^{\mathfrak{S}_{\phi}}\coloneqq[x_{i}]$ for all $i\in\{1,\dots,n\}$.
@@
It is easy to see that $\llbracket\Phi\rrbracket$ is closed under disjoint unions and that its complement is closed under homomorphisms. Corollary 3.10 implies that there exists a countable $\omega$-categorical structure with $\operatorname{CSP}(\mathfrak{B})=\llbracket\Phi\rrbracket$. $\triangle$
Remark 3.13.
We would like to point out that the existence of an $\omega$-categorical structure $\mathfrak{B}$ with the properties from Example 3.12 does not simply follow from the theorem of Cherlin, Shelah, and Shi [26], which states that for every finite set of finite connected structures $\mathcal{F}$ there exists an $\omega$-categorical structure $\mathfrak{C}$ such that a finite structure $\mathfrak{A}$ homomorphically maps to $\mathfrak{C}$ if and only if there is no homomorphism from a structure in $\mathcal{F}$ into $\mathfrak{A}$.
Formally, we claim that there is no finite set $\mathcal{F}$ of finite structures $\mathcal{F}$} such that a finite structure $\mathfrak{A}$ has a homomorphism to $\mathfrak{B}$ if and only if no structure from $\mathcal{F}$} homomorphically maps to $\mathfrak{A}$. To see this, first note that the following graphs $W_{i}$, for $i\geq 2$, satisfy the mentioned extension property: the vertex set of $W_{i}$ is $\{0,1,\dots,i\}$, there is an edge from $p$ to $q$, for $p,q\in\{0,\dots,i-1\}$, if $p-q=1\mod i$, and additionally we have the edges $(p,i)$ and $(i,p)$ for every $p\in\{0,\dots,i-1\}$. See Figure 2 for an illustration of $W_{6}$. Note that for every $i\geq 2$ the graph $W_{i}$ is a core, and that they are pairwise homomorphically incomparable. Moreover, every subgraph of $W_{i}$ does not satisfy the extension property. $\triangle$
@@ 3.3 Finite Unions of CSPs
In this section we prove that every class in GSO whose complement is closed under homomorphisms is a finite union of CSPs in GSO (Theorem 3.22); the statement announced at the beginning of Section 3 then follows (Corollary 3.24). Throughout this section, let $\tau$ be a relational signature and let $\mathcal{C}$} be a non-empty class of finite $\tau$-structures whose complement is closed under homomorphisms. In particular, $\mathcal{C}$} contains the structure $\mathfrak{I}$ with only one element where all relations are empty.
Definition 3.14.
@@
Note that the equivalence classes of $\sim$ are in one-to-one correspondence to the equivalence classes of $\sim_{0}^{\mathcal{C}}$. Also note that $\mathcal{C}$} is closed under disjoint unions if and only if $\sim$ has only one equivalence class.
If $\mathfrak{A}\in{\mathcal{C}}$, then we write $\mathfrak{A}]$ for the equivalence class of $\mathfrak{A}$ with respect to $\sim$. The following observations are immediate consequences from the definitions:
@@
    each $\sim$-equivalence class is closed under homomorphic equivalence, [^5] Two structures are said to be homomorphically equivalent if there are homomorphisms from $\mathfrak{A}$ to $\mathfrak{B}$ and from $\mathfrak{B}$ to $\mathfrak{A}$.
@@
Let $\mathfrak{A}\in{\mathcal{C}}$ and let $\mathcal{D}$ be the smallest subclass of $\mathcal{C}$ that contains $\mathfrak{A}]$ and whose complement is closed under homomorphisms. Then
@@
$\Leftrightarrow\;$ $\mathfrak{B}\uplus(\mathfrak{C}\uplus\mathfrak{D})\in{\mathcal{C}}$
$\Leftrightarrow\;$ $\mathfrak{B}^{\prime}\uplus(\mathfrak{C}\uplus\mathfrak{D})\in{\mathcal{C}}$ (since $\mathfrak{B}\sim\mathfrak{B}^{\prime}$)
$\Leftrightarrow\;$ $(\mathfrak{B}^{\prime}\uplus\mathfrak{C})\uplus\mathfrak{D}\in{\mathcal{C}}$
@@
To prove the second statement, first observe that the statement is clear if $\mathfrak{A}\in[\mathfrak{I}]$, since the complement of $\mathfrak{I}$] is closed under homomorphisms. The statement therefore follows from the assumption that $\sim$ has more than one equivalence class. Otherwise, if $\mathfrak{A}\notin[\mathfrak{I}]$, then there exists a structure $\mathfrak{B}\in{\mathcal{C}}$ such that $\mathfrak{A}\uplus\mathfrak{B}\notin{\mathcal{C}}$. Then $\mathfrak{B}\in\mathcal{C}\setminus{\mathcal{D}}$ can be shown indirectly as follows: otherwise $\mathfrak{B}$ would have a homomorphism to a structure $\mathfrak{A}^{\prime}\in[\mathfrak{A}]$. Since $\mathfrak{B}\uplus\mathfrak{A}^{\prime}$ is homomorphically equivalent to $\mathfrak{A}^{\prime}$, we have $\mathfrak{B}\uplus\mathfrak{A}^{\prime}\sim\mathfrak{A}^{\prime}\sim\mathfrak{A}$ and in particular $\mathfrak{B}\uplus\mathfrak{A}^{\prime}\in{\mathcal{C}}$. But $\mathfrak{B}\uplus\mathfrak{A}^{\prime}\in{\mathcal{C}}$ if and only if $\mathfrak{B}\uplus\mathfrak{A}\in{\mathcal{C}}$ since $\mathfrak{A}\sim\mathfrak{A}^{\prime}$. This is in contradiction to our assumption on $\mathfrak{B}$.
@@
$\triangle$
Recall that Lemma 3.5 asserts that the equivalence relation $\equiv^{\operatorname{GSO}}_{q}$ on the class of finite $\tau$-structures has finitely many equivalence classes $\mathcal{C}_{1},\dots,\mathcal{C}_{m}$, and that each of the equivalence classes $\mathcal{C}_{i}$ can be defined by a single GSO $\tau$-sentence $\Psi_{i}$ with quantifier rank $q$; we write $T^{\tau}_{q}\coloneqq\{\Psi_{1},\dots,\Psi_{m}\}$ for this set of GSO sentences.
Definition 3.18.
Let $\Phi$ be a GSO $\tau$-sentence of quantifier rank $q$. Let $J\subseteq\{1,\ldots,m\}$ be such that $\{\Psi_{j}\in T^{\tau}_{q}\mid j\in J\}$ is exactly the set of all sentences in $T^{\tau}_{q}$ that imply $\Phi$. Then $|J|$ is called the degree of $\Phi$.
It is easy to see that the degree of $\Phi$ is exactly the index of $\equiv^{\operatorname{GSO}}_{q}$ restricted to $\llbracket\Phi\rrbracket$.
@@
Let $\Phi$ be a GSO $\tau$-sentence of quantifier rank $q$ such that the complement of ${\mathcal{C}}\coloneqq\llbracket\Phi\rrbracket$ is closed under homomorphisms. Let $\sim$ be the equivalence relation from Definition 3.14 for $\mathcal{C}$}. Then for every $\sim$-class $\mathcal{D}$ there exists $I\subseteq\{1,\ldots,m\}$ such that ${\mathcal{D}}=\bigcup_{i\in I}\llbracket\Psi_{i}\rrbracket$.
Proof 3.20.
@@
From this lemma we get immediately the following corollary.